OpenStudy (anonymous):

What quadrant do the solutions lie in? 1) 3x – 2y = 10 5x + 3y = 4 A) I 2) 3x + 2y = 14 x – 3y = –10 B) II 3) x + 2y = 8 4x + 6y =22 C) III 4) 2x + 3y = –12 4x – 2y = –2 D) IV

OpenStudy (anonymous):

graph the pair of lines and see where the intersection lies in

OpenStudy (anonymous):

I have trouble solving these equations.

OpenStudy (anonymous):

Systems ... do you know how to solve systems of EQ?

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

are you allowed to use a graphing calc at school?

OpenStudy (anonymous):

I do online school.

OpenStudy (anonymous):

would you like me to show you how to solve systems of equations?

OpenStudy (anonymous):

1) 3x – 2y = 10 5x + 3y = 4 -> (2, -2) in the fourth quadrant

OpenStudy (anonymous):

I want to know the main key point in knowing how to solve problems like this.

OpenStudy (anonymous):

points*

OpenStudy (anonymous):

you're looking for the intersection of the two lines ... to find this you can use Substitution Method, Elimination Method, Graphing Method or Matrices

OpenStudy (anonymous):

Okay, I've herd of Substitution Method. What do you do for this step?

OpenStudy (anonymous):

Solve for one of the variables ... y = .....

OpenStudy (anonymous):

2) 3x + 2y = 14 x – 3y = –10 -> (2, 4) in the first quadrant!

OpenStudy (anonymous):

and plug it in the second equation and solve for x

OpenStudy (anonymous):

try solving for y in the first eq 1st system

OpenStudy (anonymous):

What do you plug in?

OpenStudy (anonymous):

Random numbers?

OpenStudy (anonymous):

did you solve for y?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

let's do that first:

OpenStudy (anonymous):

3) x + 2y = 8 4x + 6y =22 -> ( -2, 5) in the second quadrant

OpenStudy (anonymous):

do you know how to solve for y?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

ok ... here we go :

OpenStudy (anonymous):

3x-2y =10

OpenStudy (anonymous):

subtract 3x from both sides

OpenStudy (anonymous):

-2y = 10-3x

OpenStudy (anonymous):

divide by -2

OpenStudy (anonymous):

y= -5 +3/2x

OpenStudy (anonymous):

4) 2x + 3y = –12 4x – 2y = –2 -> ( -15/8, -11/4) in the third quadrant!

OpenStudy (anonymous):

this is solving for y... now we have an expresion for y we can substitute y in the second eq with the expression for y we just found

OpenStudy (anonymous):

5x+3y=4 substitute the y with -5+3/2x

OpenStudy (anonymous):

we get : 5x + 3(-5+3/2 x) = 4 ...

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

5x-15 +(9/2)x=4

OpenStudy (anonymous):

Here I like the same denominator ...(10/2)x -30/2 +(9/2)x=8/2

OpenStudy (anonymous):

now eliminate the denominator

OpenStudy (anonymous):

10x-30+9x=8

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

19x =38 -> x=2

OpenStudy (anonymous):

substitute 2 in y=-5+(3/2)x , y=-5+3 =-2 the solution (2,-2)...(IV quad)

OpenStudy (anonymous):

watch the video ...I think you'll understand better

OpenStudy (anonymous):

also use iTunes U for algebra videos ...awesome resource

OpenStudy (anonymous):

are you still here ?

OpenStudy (anonymous):

Yeah I'm here :-)

OpenStudy (anonymous):

Thanks for helping me understand this a little bit better.

OpenStudy (anonymous):

good ... did you understand ?

OpenStudy (anonymous):

use khanacademy ... Sal is good at explaining

OpenStudy (anonymous):

Lol okay :-)

OpenStudy (anonymous):

have fun ... <3Math